# What makes unlearning hard and what to do about it

### v1

@@ Abstract
Machine unlearning is the problem of removing the effect of a subset of training data (the “forget set”) from a trained model e.g. to comply with users’ requests to delete their data, or remove mislabeled, poisoned or otherwise problematic data. With unlearning research still being at its infancy, many fundamental open questions exist: Are there interpretable characteristics of forget sets that substantially affect the difficulty of the problem? How do these characteristics affect different state-of-the-art algorithms? We present the first investigation into these questions. We identify two key factors affecting unlearning difficulty and the performance of unlearning algorithms. Our evaluation on forget sets that isolate these identified factors reveals previously-unknown behaviours of state-of-the-art algorithms that don’t materialize on random forget sets. Based on our insights, we develop a framework coined Refined-Unlearning Meta-algorithm (RUM) that encompasses: (i) refining the forget set into homogenized subsets, according to different characteristics; and (ii) a meta-algorithm that employs existing algorithms to unlearn each subset and finally delivers a model that has unlearned the overall forget set. RUM substantially improves top-performing unlearning algorithms. Overall, we view our work as an important step in deepening our scientific understanding of unlearning and revealing new pathways to improving the state-of-the-art.
**footnotetext: Equal senior contribution
@@ 1 Introduction
Deep learning models have generated impressive success stories recently by leveraging increasingly large and data-hungry neural networks that are also increasingly expensive to train. This trend has led to reusing previously-trained models for a wide range of tasks more than ever before. However, the heavy reliance of deep models on training data, together with the difficulty of removing data from trained models after-the-fact, has exacerbated concerns on perpetuating harmful or outdated information, violating user privacy and other issues. Specifically, deep networks are highly non-convex, making it difficult to trace (and thus attempt to remove) the effect of a given subset of training data on the model weights. We are therefore faced with important technical challenges when it comes to building machine learning pipelines that are performant while efficiently supporting deletion requests. Machine unlearning (Nguyen et al., 2022) is a growing field that aims to address this important issue.
While unlearning is receiving increasing attention (Triantafillou et al., 2023), it is still a young area of research and the factors affecting the success of different approaches remain poorly-understood. Understanding what makes an unlearning problem easy or hard is crucial for several reasons. First, knowledge of behaviours of unlearning algorithms on different types of forgetting requests may inform which unlearning method to choose for a given request. In fact, for some requests it may be that all current methods are inadequate, suggesting that one should pay the cost of retraining from scratch rather than opting for “approximate unlearning” that imperfectly removes information after-the-fact. Further, deepening our understanding of unlearning can illuminate pathways for improving both unlearning algorithms as well as evaluation protocols by focusing on relevant factors that affect difficulty.
@@ 2.1 Unlearning problem formulation
The following definition, borrowed from (Triantafillou et al., 2023) (and similar to (Ginart et al., 2019)) formalizes this idea in a framework inspired by differential privacy (Dwork, 2006).
@@
According to the above definition, an unlearning algorithm is said to be exact unlearning if it satisfies the above definition for $\epsilon=\delta=0$, i.e., it yields a distribution of models identical to that of retraining from scratch. For neural networks, the only known exact solutions involve retraining, either naively, or in the context of mixtures where one can retrain only a subset of models affected by the deletion request (Bourtoule et al., 2021). These approaches unfortunately are inefficient; in the worst-case, even clever schemes suffer inefficiency similar to naive retraining, and may also yield poorer performance. To address this, a plethora of approximate unlearning algorithms have been recently proposed, whose $\epsilon$ and $\delta$ values aren’t known in general, but are substantially more efficient and may have higher utility.
@@
Utility and efficiency are straightforward to measure, and, in the context of classifiers, can be represented by the accuracy on the retain and test sets, and time in seconds, respectively. Measuring forgetting quality, on the other hand, is more complex and several proxies have been proposed. The simplest one is to inspect the accuracy on the forget set, with the goal of matching the accuracy on the forget set that would have been obtained by retraining from scratch. Alternatively, inspired from privacy literature Carlini et al. (2022a); Mattern et al. (2023), Membership Inference Attacks (MIAs) have been adopted by the unlearning community Kurmanji et al. (2024); Liu et al. (2024); Hayes et al. (2024) to measure forgetting quality. In essence, an MIA is designed to infer from the model’s characteristics (e.g. loss, confidence) whether a data point has been used in training, and then unlearned, versus was never trained on in the first place. Intuitively, the failure of an attacker to tell apart unlearned examples from never-seen examples marks a success for the unlearning algorithm in terms of this metric. We will consider both of these proxies in our experimental investigation.
To holistically evaluate an unlearning algorithm, we desire a single metric that captures both forgetting quality and utility. We will later introduce a “tug-of-war” metric for this purpose, inspired by Triantafillou et al. (2023).
@@ 2.2 Memorization
Deep neural networks are known to “memorize” (a subset of) their training data, with a recent theory showing that label memorization is in fact necessary for achieving close-to-optimal generalization error in classifiers (Feldman, 2020) when the data distribution is long-tailed.
@@
Memorization score (Feldman, 2020). The memorization score for an example $i\in\mathcal{D}$, with respect to a training dataset $\mathcal{D}$ and training algorithm $\mathcal{A}$ is
@@
The first term in the above equation considers models trained on all of $\mathcal{D}$ whereas the second term considers models trained on $\mathcal{D}$ excluding example $i$. Intuitively, the memorization score for an example $i$ is high if including it in training yields a different distributon of predictions on that example than excluding it from training would have. Recent works (Feldman, 2020; Feldman and Zhang, 2020; Jiang et al., 2020) identify atypical examples or outliers of the data distribution as examples that are more highly memorized: if an example has a noisy or incorrect label, the model is required to memorize it in order to predict it correctly.
@@ 3 Related Work
Fine-tune Warnecke et al. (2021); Golatkar et al. (2020) relies on catastrophic forgetting to diminish the confidence of the original model $\theta^{o}$ on $\mathcal{S}$. Catastrophic forgetting is induced by simply fine-tuning on the retain set $\mathcal{D}_{train}\setminus\mathcal{S}$. On the other hand, NegGrad Golatkar et al. (2020); Graves et al. (2021); Thudi et al. (2022) instead directly maximizes the loss on $\mathcal{S}$. This approach has been found empirically to cause a large drop in the utility of the model. To address this, NegGrad+ Kurmanji et al. (2024) combines fine-tuning and gradient ascent, by jointly minimizing the loss function on the retain set, and maximizing the loss function with respect to the forget set. SCRUB, proposed by the same authors as NegGrad+, extends the contrastive learning behind NegGrad+ by framing it as a student-teacher problem. Concretely, SCRUB is a bi-optimization algorithm, where the student aims to mimic the teacher’s behaviour on $\mathcal{R}$ and to disobey the teacher’s output with respect to $\mathcal{S}$. L1-sparse Liu et al. (2024) infuses weight “sparsity” into the unlearning algorithm by fine-tuning on the retain set with an L1-penalty, drawing inspiration from the model pruning literature Frankle and Carbin (2018); Ma et al. (2021). Influence Unlearning Izzo et al. (2021); Koh and Liang (2017) arrives at the important model’s weights by estimating how removing a data point affects $\theta^{o}$ via influence functions Cook and Weisberg (1982), and draws connections to $(\epsilon,\delta)$-forgetting Wang et al. (2022); Guo et al. (2019).
A different line of work is “relabelling-based” methods that trick the model to learn new labels for $S$. This can be achieved by finetuning the model with respect to a dataset $\mathcal{D}_{relabel}=(X_{\mathcal{S}},Y)$, where $X_{\mathcal{S}}$ are the features and labels $Y$ are sampled from a prior distribution of the label space. Saliency Unlearning (SalUn) Fan et al. (2023) learns $\mathcal{D}_{relabel}$ by optimising only the salient parameters of the model. Concretely, the authors argue that the model’s weights $\theta^{o}$ can be decomposed into salient weights and “intact” model weights, by investigating the weight space with respect to the forget set $S$ ala Smilkov et al. (2017); Adebayo et al. (2018).
@@
The closest research to ours is the contemporaneous work of Fan et al. (2024), where the authors study adversarial unlearning cases, i.e. “worst-case” forget sets. Fan et al. (2024) arrives at difficult forget sets by solving a bi-level optimization based on fine-tuning (i.e. catastrophic forgetting). Instead, we arrive at difficult partitions through the lens of interpretable factors: the degree of entanglement between the retain and forget set and memorization. While the primary aim of Fan et al. (2024) is to construct more pessimistic evaluation benchmarks, our primary aim is to deepen our understanding of unlearning problems and of the behaviour of state-of-the-art algorithms when operating on forget set of different identified characteristics, ultimately improving unlearning pipelines.
@@
(Jagielski et al., 2022) and (Toneva et al., 2018) study catastrophic forgetting during training. (Jagielski et al., 2022) finds that, when training on large datasets, examples that were only seen early in training may enjoy better privacy, in terms of MIAs and extraction attacks, compared to examples seen recently. (Toneva et al., 2018) investigate “forgetting events”, where an example that was previously correctly predicted becomes incorrectly predicted later in training. They find that examples with noisy labels witness a larger number of these forgetting events. Further, Carlini et al. (2019) find that models trained with Differential Privacy (DP) find it primarily hard to correctly predict atypical examples. We build on this literature by studying the difficulty of unlearning after-the-fact, rather than (passive) forgetting during training and draw connections to memorization, a notion closely related to atypicality in the data distribution.
@@ 4 What Makes Unlearning Hard?
In this section, we identify and empirically examine two factors that affect the difficulty of unlearning. Before diving in, we first define a simple proxy for unlearning difficulty that we will use in this section. Our goal is to capture the difficulty of performing the “balancing act” of forgetting $\mathcal{S}$ while retaining the ability to perform well on $\mathcal{R}$ and generalize well to the test set. We propose a metric to capture this “tug-of-war” (ToW) using the relative difference between the accuracies of the unlearned and the retrained model on the forget, retain and test sets, in a manner inspired by (Triantafillou et al., 2023).
@@ 4.1 The more entangled the forget and retain sets are, the harder unlearning becomes
Prior research (e.g., by Feldman et al Feldman (2020) and Carlini et al Carlini et al. (2019)) has tried to identify prototypical (or atypical) examples and their impact on learning. Primarily, this depended on the position of examples within the overall data space distribution. In contrast, here we focus on the embedding space, as unlearning depends heavily on how the model has learned to represent training data. Furthermore, instead of looking at isolated examples, we delve into how “entangled” the retain and forget sets are in embedding space. We hypothesize that higher “entanglement” leads to harder unlearning: if the two sets are highly entangled, attempting to erase $\mathcal{S}$ will cause accidentally erasing $\mathcal{R}$ too.
We propose to measure entanglement between the retain and forget sets via the below Entanglement Score (ES), inspired by a measure previously introduced in (Goldblum et al., 2020) to study learned representations.
@@
Our investigation hinges on generating three different forget/retain partitions with different degrees of entanglement: low, medium, and high. But, Equation (2) does not directly suggest a procedure for generating retain/forget partitions with a desired ES score. Hence, we achieved this indirectly using a proxy. Specifically, let $d(i,\mu;\theta^{o})=||\phi_{i}-\mu||^{2}$ denote the l2-distance in the original model’s embedding space between example $i$ and centroid $\mu$, as defined above. We compute this distance for each example in $\mathcal{D}_{train}$ and sort those examples according to their $d$-values. We then form each forget set to contain a contiguous subset of examples from different ranges of that sorted list. We find that this procedure allows us to construct retain/forget partitions of varying ES values. The ES values for our low, medium and high partitions are 309.94$\pm$98.56, 1076.99$\pm$78.64, 1612.21$\pm$110.82 for CIFAR-10, and 963.82$\pm$113.53, 2831.24$\pm$558.63, and 3876.90$\pm$426.92 for CIFAR-100. We include details of this procedure in the Section A.3, and visualizations to confirm the degree of retain/forget entanglement is in line with the scores we compute. We experiment on two datasets and two architectures, CIFAR-10/CIFAR-100 and ResNet-18/ResNet-50 with $|S|=3000$. Refer to Section A.2 for implementation details and Section A.6 for more detailed results for both datasets.
@@ 4.2 The more memorized the forget examples are, the harder unlearning becomes
Feldman et al Feldman (2020) have already established that models must memorize some atypical examples in order to perform well. Further, prior literature has also established that noisy examples (that are more likely to be memorized) witness more “forgetting events” during training (their predicted label flips to an incorrect one) Toneva et al. (2018) and that models trained with Differential Privacy (DP), a procedure where noise is added to the gradients (making it harder to memorize), find it primarily hard to correctly predict atypical examples Carlini et al. (2019). In this section, we build upon these prior insights by investigating the connection between the degree of memorization of the forget set and difficulty of unlearning.
@@
We first emphasize two key sets of conclusions. First, in terms of ToW, Figure 1(b) shows that, indeed, for most algorithms, the lower the memorization level of the forget set, the easier the problem is. In line with our prior discussion, even the original model performs well on “low-mem”, but performs very poorly on “high-mem”. Interestingly though, the two relabelling-based algorithms (SalUn and Random-label) follow an inverse trend: they perform better for higher-memorized forget sets. Second, breaking down ToW into its parts, we find interesting trends in terms of the forget set accuracy, in Figure 8. Specifically, for several unlearning algorithms, the forget accuracy for “low-mem” is still very high after unlearning them, as the model can infer the correct labels for such examples even when they weren’t included in training; this follows directly by Definition 2.2 if unlearning is done by retraining, and is shown here for the first time for approximate unlearning algorithms. On the other hand, we find that, for “high-mem”, different unlearning algorithms can (to varying degrees) cause the forget set accuracy to drop substantially; this is consistent with both Toneva et al. (2018) and Carlini et al. (2019), but shown here for the first time for approximate unlearning algorithms. Notably, we find that relabelling-based algorithms cause a larger drop in the accuracy of the forget set, relative to other approaches. This benefits ToW in the case of “high-mem” forget sets, where retraining has poor accuracy on this set (so they get rewarded by matching it), but it hurts on “low-mem”, since it causes a large discrepancy from retraining, which has high accuracy on this set (since it makes similar predictions to the original model on this set, by definition, and the original model has high accuracy on all of $\mathcal{D}_{train}$).

### v2

@@ Abstract
Machine unlearning is the problem of removing the effect of a subset of training data (the “forget set”) from a trained model e.g. to comply with users’ requests to delete their data, or remove mislabeled, poisoned or otherwise problematic data. With unlearning research still being at its infancy, many fundamental open questions exist: Are there interpretable characteristics of forget sets that substantially affect the difficulty of the problem? How do these characteristics affect different state-of-the-art algorithms? We present the first investigation into these questions. We identify two key factors affecting unlearning difficulty and the performance of unlearning algorithms. Our evaluation on forget sets that isolate these identified factors reveals previously-unknown behaviours of state-of-the-art algorithms that don’t materialize on random forget sets. Based on our insights, we develop a framework coined Refined-Unlearning Meta-algorithm (RUM) that encompasses: (i) refining the forget set into homogenized subsets, according to different characteristics; and (ii) a meta-algorithm that employs existing algorithms to unlearn each subset and finally delivers a model that has unlearned the overall forget set. RUM substantially improves top-performing unlearning algorithms. Overall, we view our work as an important step in deepening our scientific understanding of unlearning and revealing new pathways to improving the state-of-the-art.  [^3] Code is available at: https://github.com/kairanzhao/RUM
@@ 1 Introduction
Deep learning models have generated impressive success stories recently by leveraging increasingly large and data-hungry neural networks that are also increasingly expensive to train. This trend has led to reusing previously-trained models for a wide range of tasks more than ever before. However, the heavy reliance of deep models on training data, together with the difficulty of removing data from trained models after-the-fact, has exacerbated concerns on perpetuating harmful or outdated information, violating user privacy and other issues. Specifically, deep networks are highly non-convex, making it difficult to trace (and thus attempt to remove) the effect of a given subset of training data on the model weights. We are therefore faced with important technical challenges when it comes to building machine learning pipelines that are performant while efficiently supporting deletion requests. Machine unlearning [29] is a growing field that aims to address this important issue.
While unlearning is receiving increasing attention [34, 33], it is still a young area of research and the factors affecting the success of different approaches remain poorly-understood. Understanding what makes an unlearning problem easy or hard is crucial for several reasons. First, knowledge of behaviours of unlearning algorithms on different types of forgetting requests may inform which unlearning method to choose for a given request. In fact, for some requests it may be that all current methods are inadequate, suggesting that one should pay the cost of retraining from scratch rather than opting for “approximate unlearning” that imperfectly removes information after-the-fact. Further, deepening our understanding of unlearning can illuminate pathways for improving both unlearning algorithms as well as evaluation protocols by focusing on relevant factors that affect difficulty.
@@ 2.1 Unlearning problem formulation
The following definition, borrowed from [33] (and similar to [14]) formalizes this idea in a framework inspired by differential privacy [8].
@@
According to the above definition, an unlearning algorithm is said to be exact unlearning if it satisfies the above definition for $\epsilon=\delta=0$, i.e., it yields a distribution of models identical to that of retraining from scratch. For neural networks, the only known exact solutions involve retraining, either naively, or in the context of mixtures where one can retrain only a subset of models affected by the deletion request [3]. These approaches unfortunately are inefficient; in the worst-case, even clever schemes suffer inefficiency similar to naive retraining, and may also yield poorer performance. To address this, a plethora of approximate unlearning algorithms have been recently proposed, whose $\epsilon$ and $\delta$ values aren’t known in general, but are substantially more efficient and may have higher utility.
@@
Utility and efficiency are straightforward to measure, and, in the context of classifiers, can be represented by the accuracy on the retain and test sets, and time in seconds, respectively. Measuring forgetting quality, on the other hand, is more complex and several proxies have been proposed. The simplest one is to inspect the accuracy on the forget set, with the goal of matching the accuracy on the forget set that would have been obtained by retraining from scratch. Alternatively, inspired from privacy literature [4, 28], Membership Inference Attacks (MIAs) have been adopted by the unlearning community [25, 26, 20] to measure forgetting quality. In essence, an MIA is designed to infer from the model’s characteristics (e.g. loss, confidence) whether a data point has been used in training, and then unlearned, versus was never trained on in the first place. Intuitively, the failure of an attacker to tell apart unlearned examples from never-seen examples marks a success for the unlearning algorithm in terms of this metric. We will consider both of these proxies in our experimental investigation.
To holistically evaluate an unlearning algorithm, we desire a single metric that captures both forgetting quality and utility. We will later introduce a “tug-of-war” metric for this purpose, inspired by [33].
@@ 2.2 Memorization
Deep neural networks are known to “memorize” (a subset of) their training data, with a recent theory showing that label memorization is in fact necessary for achieving close-to-optimal generalization error in classifiers [11] when the data distribution is long-tailed.
@@
Memorization score [11]. The memorization score for an example $i\in\mathcal{D}$, with respect to a training dataset $\mathcal{D}$ and training algorithm $\mathcal{A}$ is
@@
The first term in the above equation considers models trained on all of $\mathcal{D}$ whereas the second term considers models trained on $\mathcal{D}$ excluding example $i$. Intuitively, the memorization score for an example $i$ is high if including it in training yields a different distribution of predictions on that example than excluding it from training would have. Recent works [11, 12, 23] identify atypical examples or outliers of the data distribution as examples that are more highly memorized: if an example has a noisy or incorrect label, the model is required to memorize it in order to predict it correctly.
@@ 3 Related Work
Fine-tune [36, 15] relies on catastrophic forgetting to diminish the confidence of the original model $\theta^{o}$ on $\mathcal{S}$. Catastrophic forgetting is induced by simply fine-tuning on the retain set $\mathcal{D}_{train}\setminus\mathcal{S}$. On the other hand, NegGrad [15, 17, 31] instead directly maximizes the loss on $\mathcal{S}$. This approach has been found empirically to cause a large drop in the utility of the model. To address this, NegGrad+ [25] combines fine-tuning and gradient ascent, by jointly minimizing the loss function on the retain set, and maximizing the loss function with respect to the forget set. SCRUB, proposed by the same authors as NegGrad+, extends the contrastive learning behind NegGrad+ by framing it as a student-teacher problem. Concretely, SCRUB is a bi-optimization algorithm, where the student aims to mimic the teacher’s behaviour on $\mathcal{R}$ and to disobey the teacher’s output with respect to $\mathcal{S}$. L1-sparse [26] infuses weight “sparsity” into the unlearning algorithm by fine-tuning on the retain set with an L1-penalty, drawing inspiration from the model pruning literature [13, 27]. Influence Unlearning [21, 24] arrives at the important model’s weights by estimating how removing a data point affects $\theta^{o}$ via influence functions [7], and draws connections to $(\epsilon,\delta)$-forgetting [35, 19].
A different line of work is “relabelling-based” methods that trick the model to learn new labels for $S$. This can be achieved by finetuning the model with respect to a dataset $\mathcal{D}_{relabel}=(X_{\mathcal{S}},Y)$, where $X_{\mathcal{S}}$ are the features and labels $Y$ are sampled from a prior distribution of the label space. Saliency Unlearning (SalUn) [10] learns $\mathcal{D}_{relabel}$ by optimising only the salient parameters of the model. Concretely, the authors argue that the model’s weights $\theta^{o}$ can be decomposed into salient weights and “intact” model weights, by investigating the weight space with respect to the forget set $S$ ala [30, 1].
@@
The closest research to ours is the contemporaneous work of [9], where the authors study adversarial unlearning cases, i.e. “worst-case” forget sets. [9] arrives at difficult forget sets by solving a bi-level optimization based on fine-tuning (i.e. catastrophic forgetting). Instead, we arrive at difficult partitions through the lens of interpretable factors: the degree of entanglement between the retain and forget set and memorization. While the primary aim of [9] is to construct more pessimistic evaluation benchmarks, our primary aim is to deepen our understanding of unlearning problems and of the behaviour of state-of-the-art algorithms when operating on forget set of different identified characteristics, ultimately improving unlearning pipelines.
@@
[22] and [32] study catastrophic forgetting during training. [22] finds that, when training on large datasets, examples that were only seen early in training may enjoy better privacy, in terms of MIAs and extraction attacks, compared to examples seen recently. [32] investigate “forgetting events”, where an example that was previously correctly predicted becomes incorrectly predicted later in training. They find that examples with noisy labels witness a larger number of these forgetting events. Further, [5] find that models trained with Differential Privacy (DP) find it primarily hard to correctly predict atypical examples. We build on this literature by studying the difficulty of unlearning after-the-fact, rather than (passive) forgetting during training and draw connections to memorization, a notion closely related to atypicality in the data distribution.
@@ 4 What Makes Unlearning Hard?
In this section, we identify and empirically examine two factors that affect the difficulty of unlearning. Before diving in, we first define a simple proxy for unlearning difficulty that we will use in this section. Our goal is to capture the difficulty of performing the “balancing act” of forgetting $\mathcal{S}$ while retaining the ability to perform well on $\mathcal{R}$ and generalize well to the test set. We propose a metric to capture this “tug-of-war” (ToW) using the relative difference between the accuracies of the unlearned and the retrained model on the forget, retain and test sets, in a manner inspired by [33].
@@ 4.1 The more entangled the forget and retain sets are, the harder unlearning becomes
Prior research (e.g., by Feldman et al [11] and Carlini et al [5]) has tried to identify prototypical (or atypical) examples and their impact on learning. Primarily, this depended on the position of examples within the overall data space distribution. In contrast, here we focus on the embedding space, as unlearning depends heavily on how the model has learned to represent training data. Furthermore, instead of looking at isolated examples, we delve into how “entangled” the retain and forget sets are in embedding space. We hypothesize that higher “entanglement” leads to harder unlearning: if the two sets are highly entangled, attempting to erase $\mathcal{S}$ will cause accidentally erasing $\mathcal{R}$ too.
We propose to measure entanglement between the retain and forget sets via the below Entanglement Score (ES), inspired by a measure previously introduced in [16] to study learned representations.
@@
Our investigation hinges on generating three different forget/retain partitions with different degrees of entanglement: low, medium, and high. But, Equation (2) does not directly suggest a procedure for generating retain/forget partitions with a desired ES score. Hence, we achieved this indirectly using a proxy. Specifically, let $d(i,\mu;\theta^{o})=||\phi_{i}-\mu||^{2}$ denote the l2-distance in the original model’s embedding space between example $i$ and centroid $\mu$, as defined above. We compute this distance for each example in $\mathcal{D}_{train}$ and sort those examples according to their $d$-values. We then form each forget set to contain a contiguous subset of examples from different ranges of that sorted list. We find that this procedure allows us to construct retain/forget partitions of varying ES values. The ES values for our low, medium and high partitions are 309.94$\pm$98.56, 1076.99$\pm$78.64, 1612.21$\pm$110.82 for CIFAR-10, and 963.82$\pm$113.53, 2831.24$\pm$558.63, and 3876.90$\pm$426.92 for CIFAR-100. We include details of this procedure in the Section A.3, along with visualizations and Maximum Mean Discrepancy (MMD) analysis to further validate ES, confirming that the degree of retain/forget entanglement aligns with the computed scores. We experiment with four dataset/architecture settings: CIFAR-10/ResNet-18, CIFAR-100/ResNet-50, Tiny-ImageNet/ResNet-18 and Tiny-ImageNet/VGG-16, using $|S|=3000$. Refer to Section A.2 for implementation details and Section A.8 for more detailed results on all datasets.
@@ 4.2 The more memorized the forget examples are, the harder unlearning becomes
Feldman et al [11] have already established that models must memorize some atypical examples in order to perform well. Further, prior literature has also established that noisy examples (that are more likely to be memorized) witness more “forgetting events” during training (their predicted label flips to an incorrect one) [32] and that models trained with Differential Privacy (DP), a procedure where noise is added to the gradients (making it harder to memorize), find it primarily hard to correctly predict atypical examples [5]. In this section, we build upon these prior insights by investigating the connection between the degree of memorization of the forget set and difficulty of unlearning.
@@
We first emphasize two key sets of conclusions. First, in terms of ToW, Figure 1(b) shows that, indeed, for most algorithms, the lower the memorization level of the forget set, the easier the problem is. In line with our prior discussion, even the original model performs well on “low-mem”, but performs very poorly on “high-mem”. Interestingly though, the two relabelling-based algorithms (SalUn and Random-label) follow an inverse trend: they perform better for higher-memorized forget sets. Second, breaking down ToW into its parts, we find interesting trends in terms of the forget set accuracy, in Figure 12. Specifically, for several unlearning algorithms, the forget accuracy for “low-mem” is still very high after unlearning them, as the model can infer the correct labels for such examples even when they weren’t included in training; this follows directly by Definition 2.2 if unlearning is done by retraining, and is shown here for the first time for approximate unlearning algorithms. On the other hand, we find that, for “high-mem”, different unlearning algorithms can (to varying degrees) cause the forget set accuracy to drop substantially; this is consistent with both [32] and [5], but shown here for the first time for approximate unlearning algorithms. Notably, we find that relabelling-based algorithms cause a larger drop in the accuracy of the forget set, relative to other approaches. This benefits ToW in the case of “high-mem” forget sets, where retraining has poor accuracy on this set (so they get rewarded by matching it), but it hurts on “low-mem”, since it causes a large discrepancy from retraining, which has high accuracy on this set (since it makes similar predictions to the original model on this set, by definition, and the original model has high accuracy on all of $\mathcal{D}_{train}$).